A TCP segment consisting of 4,256 bits of data and 184 bits of TCP header is sent to the IP layer.  The IP layer uses 200 bits of header. The IP datagram is transmitted through two networks each of which uses a 20-bit header. The target network has a maximum packet size of 1,500 bits.   

How many total bits, including headers, are delivered to the network layer protocol at the destination?
Answer #1

Answer:

given data size = 4256

header size = 184

IP layer uses header = 200

total size = 4256 + 184 + 200 = 4640

on destination side max packet size = 1500

1500 - 20 = 1480 pay load

20 bits is haeder

no of packets = 4640 / 1500 = 3.093333 = 3 ( nearest numbers)

we need 3 packets.

no of bits received at destination is

4640 + 3*20 = 4700 bits

Please give thumbsup, if you like it. Thanks.